A racing car has a uniform acceleration of 4 m s - ′2. What distance will it cover in 10 s after start?
 
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Initial Velocity of the car, u=0 ms-1
Acceleration, a= 4 m s^-2
Time, t= 10 s
We know Distance, s= ut + (1/2)at^2
Therefore, Distance covered by car in 10 second= 0 × 10 + (1/2) × 4 × 102
= 0 + (1/2) × 4× 10 × 10 m
= (1/2)× 400 m
= 200 m


Thanks